# Global parameterizations of $\pi\pi$ scattering with dispersive constraints: beyond the S0 wave

### v1

@@ III.1 P-wave parameterization
Note that both $\tau(s)$ functions are just convenient auxiliary functions to define our parameterization; they are not amplitudes by themselves. The scattering amplitude is the full $t_{1}^{(1)}(s)$.
@@
Just with five $B_{i}$ we will get good quality fits. Note that $\hat{m}_{\rho}$ is merely a parameter indicating where the phase of the conformal part alone reaches $\pi/2$. However [^4] The distinction between $\hat{m}_{\rho}$ as a parameter and the true peak position $m_{\rho}^{peak}$ was not clearly discussed in [68]., due to the $1/\sigma(s)$ factor in the conformal part, and the presence of additional non-conformal factors in the full $S$-matrix, the $\rho(770)$ peak position $m_{\rho}^{peak}$ lies near but not exactly at $\hat{m}_{\rho}$. The precise value of the peak and the energy at which the phase shift reaches $\pi/2$ must be determined using the full amplitude $t^{(1)}_{1}$, not just $\tau^{(1)}_{1,\text{conf}}$.
For the full amplitude we also need $\tau^{(1)}_{1}(s)$, which, following [68] once again, we parameterize in terms of a function $\Delta^{(1)}_{1}(s)$ as follows:
@@
Note that the whole $t^{(1)}_{1}$ amplitude in Eq. (10) exhibits the appropriate kinematic behavior around the thresholds. Namely, it behaves as $k_{\pi\omega}(s)^{2}$ near the $\pi\omega$ threshold (since parity conservation requires that the $\pi$ and $\omega$ are produced in a P wave) and as $k(s)^{3}/\sigma(s)\sim k(s)^{2}$ near the $\pi\pi$ one. In addition, since $\bar{J}_{\pi\omega}(s)$ is real below the $\pi\omega$ threshold, this ensures that the entire $t^{(1)}_{1}$ is elastic for $s<s_{\pi\omega}$. Above the $\pi\omega$ threshold, $\bar{J}_{\pi\omega}(s)$ has an imaginary part, making both $\tau^{(1)}_{1}$ and the whole $t^{(1)}_{1}$ amplitude inelastic. The advantage of using the $\bar{J}$ function is that this parameterization is analytic over the whole energy range from $s=0$ up to $(1.4{\mathrm{\,Ge\kern-1.00006ptV}})^{2}$, allowing for a straightforward continuation into the complex plane—something the usual step function does not provide.
Once again, we emphasize that this is just a convenient parameterization. In practice, we have found that four parameters $K_{i}$ with $i=0,\cdots,3$, together with the conformal parameterization in Eq. (11), are sufficient to describe the phase shift and elasticity on the real axis below 1.4 ${\mathrm{\,Ge\kern-1.00006ptV}}$. Notably, $K_{3}$ is consistently much smaller than the other three parameters, and adding a fifth term does not improve the fits.
@@
where the values of $\delta^{(1)}_{1}(s_{m})$ and $\delta^{{(1)}\,\prime}_{1}(s_{m})$ in the previous two equations are determined from the parameterization below 1.4 GeV, as outlined in the previous subsection.
@@
Both $\eta^{(1)}_{1}(s_{m})$ and its derivative are obtained from the parameterization below 1.4 GeV provided in the previous subsection. Thus, at most three free parameters are required.
@@ III.1.3 P-wave Global Fit and parameters
To begin with, for our P-wave Global Fit, shown in Fig. 1, we have used as input the values of the two first threshold parameters, $a^{(1)}_{1}$ and $b^{(1)}_{1}$, obtained from the CFD in [54] and listed in Table 10 of Appendix A. In addition, we have considered the P-wave phase shift and elasticity data collected in Fig. 1. The parameters used to obtain these curves for the P wave are given in Table 1. They were obtained by fitting the data discussed above while imposing the fulfillment of forward and Roy-like dispersion relations, as detailed in the next section.
Figure 1: We show our unconstrained and dispersively constrained P-wave Global Fit I. Top: phase shift below 1 GeV. Center: phase shift above 0.9 GeV. Bottom: elasticity from 0.9 GeV. We also show the “Old Global I” and CFD results from [68] and [54], respectively, whose inelasticity opened at $K\bar{K}$ threshold (blue vertical line). We now allow the inelasticity to start from the $\pi\omega$ threshold (red vertical line). Below that energy, only $a_{1}^{(1)}$ and $b_{1}^{(1)}$ from Table 10 and the analysis of Colangelo et al. [77] are used as input for the fit. Above, we show data from Hyams et al. [20] and Protopopescu et al. [24] although the latter are not fit. The black vertical line at 1.4 GeV indicates the matching point with the high-energy parameterization. Our Global Fit I should not be extrapolated beyond 1.85 GeV.
@@
The Global Fit I elasticity below 1.15 GeV is fairly compatible with the data from [24], but even with the inelasticity opening up at the $\pi\omega$ threshold there are four data points from Hyams et al. [20] that lie about two standard deviations below the curve. These points cannot be reconciled with the data from Protopopescu et al. [24], and our Global Fit I lies somewhere between the two experiments. Overall, allowing the inelasticity to start at the $\pi\omega$ threshold yields a smaller elasticity, consistently with the whole data set from [24] and the Hyams et al. data [20] above 1.05 GeV. It is worth noticing that the constrained Global Fit I elasticity deviates significantly from the unconstrained one between 1.2 and 1.6 GeV. This is the starting energy of the purely phenomenological fit, without any dispersive constraint. Finally, the Global Fit I uncertainties are now more uniform across the whole energy region and do not grow exaggeratedly with the energy as they did before.
However, the most relevant feature of the Global Fits is that the elasticity changes curvature between 0.9 and 1.4 GeV. In the unconstrained fits there was indeed a maximum, which is softened in the constrained Global Fit. We believe this is not an artifact, but a consequence of the well-established $\rho(1450)$ resonance, although it has not been explicitly included in our parameterizations.
In Fig. 1, we also compare our results with the dispersively constrained CFD [54] and with the “Old” Global I parameterization [68], whose curves below 1 GeV are indistinguishable. Since we now use as input an updated pion vector form factor, our Global Fit I is slightly different but still quite consistent with both of them up to the inelastic threshold. From that energy up to 1.2 GeV, the new phase shift is slightly higher than before, and from 1.2 to 1.42 GeV, they are all compatible again. From 1.42 GeV we can only compare our new Global Fit I to the old one, whose qualitative behavior is similar. Finally, the elasticity of Global Fit I also departs notably from the CFD and Old Global descriptions for all energies below 1.6 GeV,
@@ III.1.4 The three P-wave Global Fits
Despite these differences, we have been able to make the three Global Fits satisfy FDRs by deforming slightly their phase shift but visibly their elasticity. Namely, in all cases, the elasticity bump around 1.4 GeV is less pronounced than in the data, and in the 1.4 to 1.6 GeV region the dispersive representation prefers a lower $\eta(s)$ than suggested by the data. However, the three of them still present a clear change in the behavior of the elasticity from convexity to concavity and back, absent in the CFD and old global fits, and most likely related to the presence of a $\rho(1450)$ resonance.
It is important to recall that, although our global parameterizations extend up to 1.8 GeV (1.85 GeV for Global Fit I), they are only dispersively constrained up to 1.6 GeV with the $F^{0+}$ and $F^{I_{t}=1}$ FDRs. In the region beyond that energy, marked with a gray background in Fig. 2, they are just mere fits to data with the parameterization used above 1.4 GeV. Note also that we only plot our Global Fits up to the round value of the energy after the last data point included in each fit (for this wave, 1.85 GeV for Global Fit I and 1.8 GeV for Global Fits II and III). Extrapolating our fits beyond that energy is meaningless and may produce oscillations that are mere artifacts.
III.2 D0-wave parameterization
This is an attractive wave, largely dominated by the well-established $f_{2}(1270)$ resonance. For Solution I, a few data points in [20] between 0.9 and 1 GeV suggest a very small inelasticity. Thus, contrary to what was done in [54], we now allow the inelasticity of the Global Fit I to open at 0.9 GeV, which is just an effective phenomenological threshold. A non-vanishing inelasticity below the $K\bar{K}$ threshold is suggested not only by Solution I but also by the fact that the $f_{2}(1270)$ resonance, which dominates this channel, lies only one and a half widths away from the $K\bar{K}$ threshold, and its branching ratio to $4\pi$ is twice as large as that to $K\bar{K}$. However, for Global Fits II and III, we will keep the parameterization elastic up to the $K\bar{K}$ threshold, $\sqrt{s}\sim 0.992\,$ GeV, since the original data analysis [26] explicitly imposed the onset of inelasticity at this point. In that reference, a K-matrix formalism was employed with only the $\pi\pi$ and $K\bar{K}$ channels, where the $K\bar{K}$ channel was used to effectively parameterize “all the inelastic channels” [26].
As usual, we will illustrate our procedure using the Global Fit I. Below 1.4 GeV, we will adopt a parameterization very similar to that used in [54], which describes the phase shift in two parts matched at $2m_{K}$. However, now we will ensure that the matching is also differentiable. Additionally, we will extend the fit up to 1.9 GeV
@@ III.2 D0-wave parameterization
When $s^{1/2}\leq 2m_{K}$, we parameterize the phase shift by:
@@
In the intermediate energy region, $2m_{K}\leq s^{1/2}\leq 1.4\,{\mathrm{\,Ge\kern-1.00006ptV}}$, we use a rather similar parameterization:
@@
This parameterization is built up so that at the resonance peak $\eta^{(0)}_{2}=1-\epsilon$. Moreover, when the branching ratio of a naive Breit-Wigner shape to a given channel is bigger than 0.5, as it is the case here for the $\pi\pi$ decay, that branching ratio can be written as $21-\epsilon/$ 2 (otherwise it would be $\epsilon/2$). With this very naive approximation, we find an $f_{2}(1270)$ branching ratio to $\pi\pi$ of $0.871\pm 0.008$. This is consistent with the value obtained from the RPP[86], namely $\Gamma_{\pi\pi}/\Gamma_{\text{tot}}=0.843^{+0.028}_{-0.010}$, with $\Gamma_{\pi\pi}$ and $\Gamma_{\text{tot}}$ the partial width to $\pi\pi$ and the total width of the $f_{2}(1270)$, respectively.
@@
The quantities $\delta^{(0)}_{2}(s_{m})$, $\delta^{(0)}_{2}\,{}^{\prime}(s_{m})$, $\eta^{(0)}_{2}(s_{m})$ and $\eta^{(0)}_{2}\,{}^{\prime}(s_{m})$ are calculated from the parameterization below 1.4 GeV.
@@ III.2.3 D0-wave Global Fit and parameters
Note we have studied the possibility of maintaining the opening of the inelasticity at the $K\bar{K}$ threshold, as in the CFD [54]. However, as we will see in Sect. IV, forward dispersion relations clearly favor its onset at 0.9 GeV. This preference is not only evident in Global Fit I, but also in Global Fits II and III, which, as discussed below, are based on solutions whose inelasticity was set to begin at the $K\bar{K}$ threshold.
Finally, for illustration, we show in Fig. 3 how a naive Breit-Wigner (BW) formula compares to our global parameterization. This is represented by a continuous green line that covers the energy range of plus or minus one width around the peak. The BW has $\sim 85\%$ elasticity at the peak, and for its parameters we have taken the central values of the $f_{2}(1270)$ average mass and width provided at the RPP [86]. We see that for the phase shift, this shape is almost identical to our parameterization between 1.1 and 1.3 GeV. In particular, in the 1.1 to 1.2 GeV region where the two data sets are in clear tension, it is very close to our parameterization, that lies in between both sets. However, starting around 1.3 GeV, there is a growing deviation of the BW with respect to our global parameterization, which is more pronounced in the elasticity. This asymmetry between the two sides of the Breit-Wigner may be due to the presence of the $K\bar{K}$ threshold on the left but also to the well-established $f_{2}^{\prime}(1525)$ resonance on the right, whose estimated width at the RPP is $86\pm 5{\mathrm{\,Me\kern-1.00006ptV}}$. Nevertheless, the $f_{2}^{\prime}(1525)$ coupling to two pions is very small, and thus, its shape might go almost unnoticed in $\pi\pi$ scattering, while still producing a small deformation in the naively expected $f_{2}(1270)$ shape.

### v2

@@ III.1 P-wave parametrization
Figure 1: We show our unconstrained and dispersively constrained P-wave Global Fit I. Top: phase shift below 1 GeV. Center: phase shift above 0.9 GeV. Bottom: elasticity from 0.9 GeV. We also show the “Old Global I” and CFD results from [68] and [54], respectively, whose inelasticity opened at $K\bar{K}$ threshold (blue vertical line). We now allow the inelasticity to start from the $\pi\omega$ threshold (red vertical line). Below that energy, only $a_{1}^{(1)}$ and $b_{1}^{(1)}$ from Table 10 and the analysis of Colangelo et al. [77] are used as input for the fit. Above, we show data from Hyams et al. [20] and Protopopescu et al. [24] although the latter are not fit. The black vertical line at 1.4 GeV indicates the matching point with the high-energy parametrization. Our Global Fit I should not be extrapolated beyond 1.85 GeV.
@@
Note that both $\tau(s)$ functions are just convenient auxiliary functions to define our parametrization; they are not amplitudes by themselves. The scattering amplitude is the full $t_{1}^{(1)}(s)$.
As we commented before, we are not describing the full $S$-matrix, just its first diagonal element $S_{11}$, and unitarity just imposes $\eta(s)=|S_{11}|\leq 1$ (for simplicity, spin and isospin indices are suppressed). For this reason, the factorization scheme $S_{11}=\Pi_{k}S_{11}^{\{k\}}$ with $|S^{\{k\}}_{11}|\leq 1$, is common in the literature [87, 88, 89, 65, 67]. In principle, below a given $k$-th threshold $S_{11}^{\{k\}}=1$ but, with an appropriate momentum dependence, it becomes less than one above. This description is enough to describe $S_{11}$ in the real axis as we do here. For other purposes, each $S_{11}^{\{k\}}$ can be given a more detailed analytic structure.
@@
Just with five $B_{i}$ we get good quality fits. Note that $\hat{m}_{\rho}$ is merely a parameter indicating where the phase of the conformal part alone reaches $\pi/2$. However [^4] The distinction between $\hat{m}_{\rho}$ as a parameter and the true peak position $m_{\rho}^{peak}$ was not clearly discussed in [68]., due to the $1/\sigma(s)$ factor in the conformal part, and the presence of additional non-conformal factors in the full $S$-matrix, the $\rho(770)$ peak position $m_{\rho}^{peak}$ lies near but not exactly at $\hat{m}_{\rho}$. The precise value of the peak and the energy at which the phase shift reaches $\pi/2$ must be determined using the full amplitude $t^{(1)}_{1}$, not just $\tau^{(1)}_{1,\text{conf}}$.
For the full amplitude we also need $\tau^{(1)}_{1}(s)$, which, following [68] once again, we parametrize in terms of a function $\Delta^{(1)}_{1}(s)$ as follows:
@@
Note that the whole $t^{(1)}_{1}$ amplitude in Eq. (10) exhibits the appropriate kinematic behavior around the thresholds. Namely, it behaves as $k_{\pi\omega}(s)^{2}$ near the $\pi\omega$ threshold (since parity conservation requires that the $\pi$ and $\omega$ are produced in a P wave) and as $k(s)^{3}/\sigma(s)\sim k(s)^{2}$ near the $\pi\pi$ one. In addition, since $\bar{J}_{\pi\omega}(s)$ is real below the $\pi\omega$ threshold, this ensures that the entire $t^{(1)}_{1}$ is elastic for $s<s_{\pi\omega}$. Above the $\pi\omega$ threshold, $\bar{J}_{\pi\omega}(s)$ has an imaginary part, making both $\tau^{(1)}_{1}$ and the whole $t^{(1)}_{1}$ amplitude inelastic. The advantage of using the $\bar{J}$ function is that this parametrization is analytic over the whole energy range from $s=0$ up to $(1.4{\mathrm{\,Ge\kern-1.00006ptV}})^{2}$, allowing for a straightforward continuation into the complex plane—something the usual step function does not provide.
Once again, we emphasize that this is just a convenient parametrization. In practice, we have found that four parameters $K_{i}$ with $i=0,\cdots,3$, together with the conformal parametrization in Eq. (11), are sufficient to describe the phase shift and elasticity on the real axis below 1.4 ${\mathrm{\,Ge\kern-1.00006ptV}}$. Notably, $K_{3}$ is consistently much smaller than the other three parameters, and adding a fifth term does not improve the fits.
@@
where the values of $\delta^{(1)}_{1}(s_{m})$ and $\delta^{{(1)}\,\prime}_{1}(s_{m})$ in the previous two equations are determined from the parametrization below 1.4 GeV, as outlined in the previous subsection.
@@
Both $\eta^{(1)}_{1}(s_{m})$ and its derivative are obtained from the parametrization below 1.4 GeV provided in the previous subsection. Thus, at most three free parameters are required.
@@ III.1.3 P-wave Global Fit and parameters
To begin with, for our P-wave Global Fit, shown in Fig. 1, we have used as input the values of the two first threshold parameters, $a^{(1)}_{1}$ and $b^{(1)}_{1}$, obtained from the CFD in [54] and listed in Table 10 of Appendix A. In addition, we have considered the P-wave phase shift and elasticity data collected in Fig. 1. The parameters used to plot these curves for the P wave are given in Table 1. They were obtained by fitting the data discussed above while imposing the fulfillment of forward and Roy-like dispersion relations, as detailed in the next section.
@@
The Global Fit I elasticity below 1.15 GeV is fairly compatible with the data from [24], but even with the inelasticity opening up at the $\pi\omega$ threshold there are four data points from Hyams et al. [20] that lie about two standard deviations below the curve. These points cannot be reconciled with the data from Protopopescu et al. [24], and our Global Fit I lies somewhere between the two experiments. Overall, allowing the inelasticity to start at the $\pi\omega$ threshold yields a smaller elasticity, consistently with the whole data set from [24] and the Hyams et al. data [20] above 1.05 GeV. It is worth noticing that the constrained Global Fit I elasticity deviates significantly from the unconstrained one between 1.2 and 1.6 GeV. This is the starting energy of the purely phenomenological fit, without any dispersive constraint. Finally, the Global Fit I uncertainties are now more uniform across the whole energy region and do not grow exaggeratedly with the energy as they did before in [54, 68].
However, the most relevant feature of the Global Fits is that the elasticity changes curvature between 0.9 and 1.4 GeV. In the unconstrained fits there was indeed a maximum, which is softened in the constrained Global Fit. We believe this is not an artifact, but a consequence of the well-established $\rho(1450)$ resonance, although it has not been explicitly included in our parametrizations.
@@
In Fig. 1, we also compare our results with the dispersively constrained CFD [54] and with the “Old” Global I parametrization [68], whose curves below 1 GeV are indistinguishable. Since we now use as input an updated pion vector form factor, our Global Fit I is slightly different but still quite consistent with both of them up to the inelastic threshold. From that energy up to 1.2 GeV, the new phase shift is slightly higher than before, and from 1.2 to 1.42 GeV, they are all compatible again. From 1.42 GeV we can only compare our new Global Fit I to the old one, whose qualitative behavior is similar. Finally, the elasticity of Global Fit I also departs notably from the CFD and Old Global descriptions for all energies below 1.6 GeV.
@@ III.1.4 The three P-wave Global Fits
Despite these differences, we have been able to make the three Global Fits satisfy FDRs by deforming slightly their phase shift but visibly their elasticity. Namely, in all cases, the elasticity bump around 1.4 GeV is less pronounced than in the data, and in the 1.4 to 1.6 GeV region the dispersive representation prefers a lower $\eta(s)$ than suggested by the data. However, the three of them still present a clear change in the behavior of the elasticity from convexity to concavity and back, absent in the CFD and old global fits, and most likely related to the presence of the $\rho(1450)$ resonance.
It is important to recall that, although our global parametrizations extend up to 1.8 GeV (1.85 GeV for Global Fit I), they are only dispersively constrained up to 1.6 GeV with the $F^{0+}$ and $F^{I_{t}=1}$ FDRs. In the region beyond that energy, marked with a gray background in Fig. 2, they are just mere fits to data with the parametrization used above 1.4 GeV. Note also that we only plot our Global Fits up to the round value of the energy after the last data point included in each fit (for this wave, 1.85 GeV for Global Fit I and 1.8 GeV for Global Fits II and III). Extrapolating our fits beyond that energy is meaningless and may produce oscillations that are mere artifacts.
III.2 D0-wave parametrization
This is an attractive wave, largely dominated by the well-established $f_{2}(1270)$ resonance. For Solution I, a few data points in [20] between 0.9 and 1 GeV suggest a very small inelasticity. Thus, contrary to what was done in [54], we now allow the inelasticity of the Global Fit I to open at 0.9 GeV, which is just an effective phenomenological threshold. A non-vanishing inelasticity below the $K\bar{K}$ threshold is suggested not only by Solution I but also by the fact that the $f_{2}(1270)$ resonance, which dominates this channel, lies only one and a half widths away from the $K\bar{K}$ threshold, and its branching ratio to $4\pi$ is twice as large as that to $K\bar{K}$. However, for Global Fits II and III, we will keep the parametrization elastic up to the $K\bar{K}$ threshold, $\sqrt{s}\sim 0.992\,$ GeV, since the original data analysis [26] explicitly imposed the onset of inelasticity at this point. In that reference, a K-matrix formalism was employed with only the $\pi\pi$ and $K\bar{K}$ channels, where the $K\bar{K}$ channel was used to effectively parametrize “all the inelastic channels” [26].
As usual, we will illustrate our procedure using the Global Fit I. Below 1.4 GeV, we will adopt a parametrization very similar to that used in [54], which describes the phase shift in two parts matched at $2m_{K}$. However, now we will ensure that the matching is also differentiable. Additionally, we will extend the fit up to 1.9 GeV.
@@ III.2 D0-wave parametrization
When $s^{1/2}\leq 2m_{K}$, we parametrize the phase shift by:
@@
In the intermediate energy region, $2m_{K}\leq s^{1/2}\leq 1.4\,{\mathrm{\,Ge\kern-1.00006ptV}}$, we use a rather similar parametrization:
@@
This parametrization is built up so that at the resonance peak $\eta^{(0)}_{2}=1-\epsilon$. Moreover, when the branching ratio of a naive Breit-Wigner shape to a given channel is bigger than 0.5, as it is the case here for the $\pi\pi$ decay, that branching ratio can be written as $21-\epsilon/$ 2 (otherwise it would be $\epsilon/2$). With this very naive approximation, we find an $f_{2}(1270)$ branching ratio to $\pi\pi$ of $0.871\pm 0.008$. This is consistent with the value obtained from the RPP[90], namely $\Gamma_{\pi\pi}/\Gamma_{\text{tot}}=0.843^{+0.028}_{-0.010}$, with $\Gamma_{\pi\pi}$ and $\Gamma_{\text{tot}}$ the partial width to $\pi\pi$ and the total width of the $f_{2}(1270)$, respectively.
@@
The quantities $\delta^{(0)}_{2}(s_{m})$, $\delta^{(0)}_{2}\,{}^{\prime}(s_{m})$, $\eta^{(0)}_{2}(s_{m})$ and $\eta^{(0)}_{2}\,{}^{\prime}(s_{m})$ are calculated from the parametrization below 1.4 GeV.
@@ III.2.3 D0-wave Global Fit and parameters
Note we have studied the possibility of maintaining the opening of the inelasticity at the $K\bar{K}$ threshold, as in the CFD [54]. However, as we will see in Sec. IV, forward dispersion relations clearly favor its onset at 0.9 GeV. This preference is not only evident in Global Fit I, but also in Global Fits II and III, which, as discussed below, are based on solutions whose inelasticity was set to begin at the $K\bar{K}$ threshold.
Finally, for illustration, we show in Fig. 3 how a naive Breit-Wigner (BW) formula compares to our global parametrization. This is represented by a continuous green line that covers the energy range of plus or minus one width around the peak. The BW has $\sim 85\%$ elasticity at the peak, and for its parameters we have taken the central values of the $f_{2}(1270)$ average mass and width provided at the RPP [90]. We see that for the phase shift, this shape is almost identical to our parametrization between 1.1 and 1.3 GeV. In particular, in the 1.1 to 1.2 GeV region where the two data sets are in clear tension, it is very close to our parametrization, that lies in between both sets. However, starting around 1.3 GeV, there is a growing deviation of the BW with respect to our global parametrization, which is more pronounced in the elasticity. This asymmetry between the two sides of the Breit-Wigner may be due to the presence of the $K\bar{K}$ threshold on the left but also to the well-established $f_{2}^{\prime}(1525)$ resonance on the right, whose estimated width at the RPP is $86\pm 5{\mathrm{\,Me\kern-1.00006ptV}}$. Nevertheless, the $f_{2}^{\prime}(1525)$ coupling to two pions is very small, and thus, its shape might go almost unnoticed in $\pi\pi$ scattering, while still producing a small deformation in the naively expected $f_{2}(1270)$ shape.